# Localized RETE for Incremental Graph Queries with Nested Graph Conditions

### v1

@@ A.5. Supplementary Lemmata
For a nested condition of the form $\psi=\exists(a:Q\rightarrow Q^{\prime},\psi^{\prime})$, $O$ is given by $O=order(N^{\Phi}_{Q})\circ toposort^{-1}(RPS_{l})\circ order^{\Psi}(N^{\Phi}_{(Q^{\prime},\psi^{\prime})})\circ[\ltimes]^{\Phi}$. By Lemma 13 in [BG24b], it follows that $\mathcal{C}^{\Phi}_{1}=execute(order(N^{\Phi}_{Q}),X^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{0})$ is consistent for all nodes in $N^{\Phi}_{Q}$. Since no node in $N^{\Phi}_{Q}$ depends on a node in $RPS_{l}$, $\mathcal{C}^{\Phi}_{2}=execute(toposort^{-1}(RPS_{l}),X^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{1})$ must then be consistent for all nodes in $N^{\Phi}_{Q}\cup RPS_{l}$. By the same argument and the induction hypothesis, $\mathcal{C}^{\Phi}_{3}=execute(order^{\Psi}(N^{\Phi}_{(Q^{\prime},\psi^{\prime})}),X^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{2})$ is consistent for all nodes in $N^{\Phi}_{Q}\cup RPS_{l}\cup N^{\Phi}_{(Q^{\prime},\psi^{\prime})}$ and $\mathcal{C}^{\Phi}_{4}=execute([\ltimes]^{\Phi},X^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{3})=execute(O,N^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{0})$ is consistent for all nodes in $N^{\Phi}$.
@@
For a nested condition of the form $\psi=\psi_{1}\wedge\psi_{2}$, $O$ is given by $order^{\Psi}(N^{\Phi}=order(N^{\Phi}_{Q})\circ toposort^{-1}(RPS_{1})\circ order^{\Psi}(N^{\Phi}_{(Q,\psi_{1})})\circ[\ltimes]^{\Phi}_{1}\circ toposort^{-1}(RPS_{2})\circ order^{\Psi}(N^{\Phi}_{(Q,\psi_{2})})\circ[\ltimes]^{\Phi}_{2}$. By the same argumentation as for the case where $\psi=\exists(a:Q\rightarrow Q^{\prime},\psi^{\prime})$, the configuration $\mathcal{C}^{\Phi}_{1}=execute(O_{1},N^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{0})$ is consistent for all nodes in $N^{\Phi}_{Q}\cup RPS_{1}\cup N^{\Phi}_{(Q,\psi_{1})}\cup[\ltimes]^{\Phi}_{1}$, where $O_{1}=order(N^{\Phi}_{Q})\circ toposort^{-1}(RPS_{1})\circ order^{\Psi}(N^{\Phi}_{(Q,\psi_{1})})\circ[\ltimes]^{\Phi}_{1}$. By the induction hypothesis and the construction of $N^{\Phi}$, it then holds that $\mathcal{C}^{\Phi}_{2}=execute(O_{2},N^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{1})=execute(O,N^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{0})$ is consistent for all nodes in $N^{\Phi}$, where $O_{2}=toposort^{-1}(RPS_{2})\circ order^{\Psi}(N^{\Phi}_{(Q,\psi_{2})})\circ[\ltimes]^{\Phi}_{2}$.
@@
Let $H$ be a graph, $H_{p}\subseteq H$, $[\ltimes]^{\Phi}\in V^{N^{\Phi}}$ a marking-sensitive semi-join node with left dependency $n^{\Phi}_{l}$ and right dependency $n^{\Phi}_{r}$, and $\mathcal{C}^{\Phi}_{0}$ a configuration such that $\mathcal{C}^{\Phi}_{0}([\ltimes]^{\Phi})=\{\}$. Executing $[\ltimes]^{\Phi}$ via $\mathcal{C}^{\Phi}_{1}=execute([\ltimes]^{\Phi},N^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{0})$ then takes $O(|\mathcal{C}^{\Phi}_{0}(n^{\Phi}_{l})|)$ steps.
@@
Let $H$ be a graph, $H_{p}\subseteq H$, $[\rhd]^{\Phi}\in V^{N^{\Phi}}$ a marking-sensitive anti-join node with left dependency $n^{\Phi}_{l}$ and right dependency $n^{\Phi}_{r}$, and $\mathcal{C}^{\Phi}_{0}$ a configuration such that $\mathcal{C}^{\Phi}_{0}([\ltimes]^{\Phi})=\{\}$. Executing $[\ltimes]^{\Phi}$ via $\mathcal{C}^{\Phi}_{1}=execute([\ltimes]^{\Phi},N^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{0})$ then takes $O(|\mathcal{C}^{\Phi}_{0}(n^{\Phi}_{l})|)$ steps.
@@
Let $H$ be an edge-dominated graph, $H_{p}\subseteq H$, $(N,p)$ a RETE net created via the procedure described in [BG23] for the extended graph query $(Q,\psi)$, $\mathcal{C}$ a consistent configuration for $(N,p)$ for host graph $H$, and $\mathcal{C}^{\Phi}$ a consistent configuration for the marking-sensitive RETE net $(N^{\Phi},p^{\Phi})=localize^{NGC}(Q,\psi)$ for host graph $H$ and relevant subgraph $H_{p}$ corresponding to $(N,p)$. It then holds that $\sum_{n^{\Phi}\in V^{N^{\Phi}}}\mathcal{C}^{\Phi}(n^{\Phi})|\leq 10\cdot\sum_{n\in V^{N}}\mathcal{C}(n)|$.
@@
In the base case where $\psi=true$, it holds that $localize^{\Psi}(Q,\psi)=localize(N,p)$. From Lemma 38, it then immediately follows that $\sum_{n^{\Phi}\in V^{N^{\Phi}}}\mathcal{C}^{\Phi}(n^{\Phi})|\leq 10\cdot\sum_{n\in V^{N}}\mathcal{C}(n)|$.
@@
For a nested condition of the form $\psi=\exists(a:Q\rightarrow Q^{\prime},\psi^{\prime})$, $N^{\Phi}$ consists of the localized RETE net for $Q$, $(N^{\Phi}_{Q},p^{\Phi}_{Q})=localize(Q)$, the localized RETE net $(N^{\Phi}_{(Q^{\prime},\psi^{\prime})},p^{\Phi}_{(Q^{\prime},\psi^{\prime})})=localize^{\Psi}(Q^{\prime},\psi^{\prime})$, a request projection structure $RPS_{l}=RPS(p^{\Phi}_{Q},N^{\Phi}_{(Q^{\prime},\psi^{\prime})})$, and a marking-sensitive semi-join node $p^{\Phi}=[\ltimes]^{\Phi}$.
@@
In the worst case, the extension points of $(N^{\Phi}_{(Q^{\prime},\psi^{\prime})},p^{\Phi}_{(Q^{\prime},\psi^{\prime})})$ already contain all possible single-vertex matches into $H$ even without the additional input via $RPS_{l}$. Therefore, by the induction hypothesis, it must hold that $\sum_{n^{\Phi}\in V^{N^{\Phi}_{(Q^{\prime},\psi^{\prime})}}}\mathcal{C}^{\Phi}(n^{\Phi})|\leq 10\cdot\sum_{n\in V^{N^{\prime}}}\mathcal{C}(n)|$.
From Lemma 12 in [BG24b], it follows that $\sum_{n^{\Phi}_{Q}\in V^{RPS_{l}}}|\mathcal{C}^{\Phi}(n^{\Phi}_{Q})|\leq 3\cdot|\mathcal{C}(p_{Q})|$.
@@
Consequently, it must hold that $\sum_{n^{\Phi}\in V^{N^{\Phi}}}\mathcal{C}^{\Phi}(n^{\Phi})|\leq 10\cdot\sum_{n\in V^{N}}\mathcal{C}(n)|$.
For a nested condition of the form $\psi=\neg\psi^{\prime}$, $N^{\Phi}$ consists of the localized RETE net for the plain pattern $Q$, $(N^{\Phi}_{Q},p^{\Phi}_{Q})=localize(Q)$, the localized RETE net $(N^{\Phi}_{(Q,\psi^{\prime})},p^{\Phi}_{(Q,\psi^{\prime})})=localize^{\Psi}(Q,\psi^{\prime})$, a request projection structure $RPS_{l}=RPS(p^{\Phi}_{Q},N^{\Phi}_{(Q,\psi^{\prime})})$, and a marking-sensitive anti-join node $p^{\Phi}=[\rhd]^{\Phi}$.
@@
By analogous argumentation as for the existential case, it must then also hold that $\sum_{n^{\Phi}\in V^{N^{\Phi}}}\mathcal{C}^{\Phi}(n^{\Phi})|\leq 10\cdot\sum_{n\in V^{N}}\mathcal{C}(n)|$.
For a nested condition of the form $\psi=\psi_{1}\wedge\psi_{2}$, $N^{\Phi}$ consists of the localized RETE net for the plain pattern $Q$, $(N^{\Phi}_{Q},p^{\Phi}_{Q})=localize(Q)$, the localized RETE net $(N^{\Phi}_{(Q,\psi_{1})},p^{\Phi}_{(Q,\psi_{1})})=localize^{\Psi}(Q,\psi_{1})$, a request projection structure $RPS_{1}=RPS(p^{\Phi}_{Q},N^{\Phi}_{(Q,\psi_{1})})$, a marking-sensitive semi-join node $[\ltimes]^{\Phi}_{1}$, the localized RETE net $(N^{\Phi}_{(Q,\psi_{2})},p^{\Phi}_{(Q,\psi_{2})})=localize^{\Psi}(Q,\psi_{2})$, a request projection structure $RPS_{2}=RPS([\ltimes]^{\Phi}_{1},N^{\Phi}_{(Q,\psi_{2})})$, and a marking-sensitive semi-join node $p^{\Phi}=[\ltimes]^{\Phi}_{2}$.
@@
As in the existential case, it holds that $\sum_{n^{\Phi}\in V^{N^{\Phi}_{Q}}}|\mathcal{C}^{\Phi}(n^{\Phi})|\leq 10\sum_{n\in V^{N_{Q}}\setminus\{p_{Q}\}}|\mathcal{C}(n)|+7\cdot|\mathcal{C}(p_{Q})|$. Furthermore, it must hold that $\sum_{n^{\Phi}\in V^{N^{\Phi}_{(Q,\psi_{1})}}}|\mathcal{C}^{\Phi}(n^{\Phi})|\leq 10\cdot\sum_{n\in V^{N_{1}}}\mathcal{C}(n)|$ and $\sum_{n^{\Phi}\in V^{N^{\Phi}_{(Q,\psi_{1})}}}|\mathcal{C}^{\Phi}(n^{\Phi})|\leq 10\cdot\sum_{n\in V^{N_{2}}}\mathcal{C}(n)|$ by the induction hypothesis.
From Lemma 12 in [BG24b], it follows that $\sum_{n^{\Phi}_{Q}\in V^{RPS_{1}}}|\mathcal{C}^{\Phi}(n^{\Phi}_{Q})|\leq 3\cdot|\mathcal{C}(p_{Q})|$.
We furthermore know by Theorem 8 that $\mathcal{C}^{\Phi}([\ltimes]^{\Phi}_{1})$ can only contain matches for $Q$ that satisfy $\phi_{1}$ and $\mathcal{C}([\ltimes]_{1})$ contains all such matches. From Lemma 12 in [BG24b], it thus follows that $|\mathcal{C}^{\Phi}([\ltimes]^{\Phi}_{1})|\leq|\mathcal{C}([\ltimes]_{1})|$. Consequently, it must hold that $\sum_{n^{\Phi}_{Q}\in V^{RPS_{2}}}|\mathcal{C}^{\Phi}(n^{\Phi}_{Q})|\leq 3\cdot|\mathcal{C}([\ltimes]_{1})|$. By Theorem 8 and Lemma 12 in [BG24b], it also follows that $|\mathcal{C}^{\Phi}([\ltimes]^{\Phi}_{2})|\leq|\mathcal{C}([\ltimes]_{2})|$.
We thus know that $\sum_{n^{\Phi}\in V^{N^{\Phi}}}\mathcal{C}^{\Phi}(n^{\Phi})|\leq 10\cdot\sum_{n\in V^{N}}\mathcal{C}(n)|$.
@@
Let $H$ be a graph, $H_{p}\subseteq H$, and $(Q,\psi)$ an extended graph query. Furthermore, let $\mathcal{C}^{\Phi}$ be a consistent configuration for the modular extension $(X^{sat},p^{sat})$ of $(N^{sat},p^{sat})=localize^{sat}(Q,\psi)$. It then holds that $\forall m\in\mathcal{M}^{Q}_{H}:m\text{ is subgraph satisfaction dependent }\Rightarrow(m,\infty)\in\mathcal{C}^{\Phi}(p^{sat})$.
@@
For a nested condition of the form $\psi=\exists(a:Q\rightarrow Q^{\prime},\psi^{\prime})$, $(N^{sat},p^{sat})=localize^{sat}(Q,\psi)$ consists of the localized RETE net for the plain pattern $Q$, $(N^{\Phi}_{Q},p^{\Phi}_{Q})=localize(Q)$, the localized RETE net for the plain pattern $Q^{\prime}$, $(N^{\Phi}_{Q^{\prime}},p^{\Phi}_{Q^{\prime}})=localize(Q^{\prime})$, the RETE net $(N^{sat}_{(Q^{\prime},\psi^{\prime})},p^{sat}_{(Q^{\prime},\psi^{\prime})})=localize^{sat}(Q^{\prime},\psi^{\prime})$, a marking-sensitive union node $[\cup]^{\Phi}$ with dependencies $p^{\Phi}_{Q^{\prime}}$ and $p^{sat}_{(Q^{\prime},\psi^{\prime})}$, a request projection structure $RPS_{r}=RPS([\cup]^{\Phi},N^{\Phi}_{Q})$, and a marking-sensitive semi-join node $p^{sat}=[\ltimes]^{\Phi}$ with left dependency $p^{\Phi}_{Q}$ and right dependency $[\cup]^{\Phi}$.
By Lemma 1 in [BG24b], the induction hypothesis, and the semantics definition of the marking-sensitive union node, it must then hold that $\forall m^{\prime}\in\mathcal{M}^{Q^{\prime}}_{H}:m^{\prime}(Q^{\prime})\cap H_{p}\neq\emptyset\vee m^{\prime}\text{ is subgraph satisfaction dependent }\Rightarrow(m^{\prime},\infty)\in\mathcal{C}^{\Phi}([\cup]^{\Phi})$.
From the construction of $RPS_{r}$ and Lemma 1 in [BG24b], it follows that $\forall m\in\mathcal{M}^{Q}_{H}:(\exists m^{\prime}\in\mathcal{M}^{Q^{\prime}}_{H}:m=m^{\prime}\circ a\wedge(m^{\prime}(Q^{\prime})\cap H_{p}\neq\emptyset\vee m^{\prime}\text{ is subgraph satisfaction dependent}))\Rightarrow(m,\infty)\in\mathcal{C}^{\Phi}(p^{\Phi}_{Q})$.
@@
Let $H$ be a graph, $H_{p}\subseteq H$, $(Q,\psi)$ a graph query, and $(X^{sat},p_{X}^{sat})$ a modular extension of the marking-sensitive RETE net $(N^{sat},p^{sat})=localize^{sat}(Q,\psi)$. Furthermore, let $\mathcal{C}^{\Phi}_{0}$ be an arbitrary configuration for $(X^{sat},p_{X}^{sat})$. Executing $(X^{sat},p_{X}^{sat})$ via $O=order^{sat}(N^{sat})$ then yields a configuration $\mathcal{C}^{\Phi}=execute(O,N^{sat},H,H_{p},\mathcal{C}^{\Phi}_{0})$ that is consistent for all nodes in $N^{sat}$.

### v2

@@ A.5. Supplementary Lemmata
For a nested condition of the form $\psi=\exists(a:Q\rightarrow Q^{\prime},\psi^{\prime})$, $O$ is given by $O=order(N^{\Phi}_{Q})\circ toposort^{-1}(RPS^{\infty}_{l})\circ order^{\Psi}(N^{\Phi}_{(Q^{\prime},\psi^{\prime})})\circ[\ltimes]^{\Phi}$. By Lemma 13 in [BG24b], it follows that $\mathcal{C}^{\Phi}_{1}=execute(order(N^{\Phi}_{Q}),X^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{0})$ is consistent for all nodes in $N^{\Phi}_{Q}$. Since no node in $N^{\Phi}_{Q}$ depends on a node in $RPS^{\infty}_{l}$, $\mathcal{C}^{\Phi}_{2}=execute(toposort^{-1}(RPS^{\infty}_{l}),X^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{1})$ must then be consistent for all nodes in $N^{\Phi}_{Q}\cup RPS^{\infty}_{l}$. By the same argument and the induction hypothesis, $\mathcal{C}^{\Phi}_{3}=execute(order^{\Psi}(N^{\Phi}_{(Q^{\prime},\psi^{\prime})}),X^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{2})$ is consistent for all nodes in $N^{\Phi}_{Q}\cup RPS^{\infty}_{l}\cup N^{\Phi}_{(Q^{\prime},\psi^{\prime})}$ and $\mathcal{C}^{\Phi}_{4}=execute([\ltimes]^{\Phi},X^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{3})=execute(O,N^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{0})$ is consistent for all nodes in $N^{\Phi}$.
@@
For a nested condition of the form $\psi=\psi_{1}\wedge\psi_{2}$, $O$ is given by $order^{\Psi}(N^{\Phi})=order(N^{\Phi}_{Q})\circ toposort^{-1}(RPS^{\infty}_{1})\circ order^{\Psi}(N^{\Phi}_{(Q,\psi_{1})})\circ[\ltimes]^{\Phi}_{1}\circ toposort^{-1}(RPS^{\infty}_{2})\circ order^{\Psi}(N^{\Phi}_{(Q,\psi_{2})})\circ[\ltimes]^{\Phi}_{2}$. By the same argumentation as for the case where $\psi=\exists(a:Q\rightarrow Q^{\prime},\psi^{\prime})$, the configuration $\mathcal{C}^{\Phi}_{1}=execute(O_{1},N^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{0})$ is consistent for all nodes in $N^{\Phi}_{Q}\cup RPS^{\infty}_{1}\cup N^{\Phi}_{(Q,\psi_{1})}\cup[\ltimes]^{\Phi}_{1}$, where $O_{1}=order(N^{\Phi}_{Q})\circ toposort^{-1}(RPS^{\infty}_{1})\circ order^{\Psi}(N^{\Phi}_{(Q,\psi_{1})})\circ[\ltimes]^{\Phi}_{1}$. By the induction hypothesis and the construction of $N^{\Phi}$, it then holds that $\mathcal{C}^{\Phi}_{2}=execute(O_{2},N^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{1})=execute(O,N^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{0})$ is consistent for all nodes in $N^{\Phi}$, where $O_{2}=toposort^{-1}(RPS^{\infty}_{2})\circ order^{\Psi}(N^{\Phi}_{(Q,\psi_{2})})\circ[\ltimes]^{\Phi}_{2}$.
@@
(Execution time of semi-join nodes is linear in the number of matches for the semi-join’s left dependency.) Let $H$ be a graph, $H_{p}\subseteq H$, $[\ltimes]^{\Phi}\in V^{N^{\Phi}}$ a marking-sensitive semi-join node with left dependency $n^{\Phi}_{l}$ and right dependency $n^{\Phi}_{r}$, and $\mathcal{C}^{\Phi}_{0}$ a configuration such that $\mathcal{C}^{\Phi}_{0}([\ltimes]^{\Phi})=\{\}$. Executing $[\ltimes]^{\Phi}$ via $\mathcal{C}^{\Phi}_{1}=execute([\ltimes]^{\Phi},N^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{0})$ then takes $O(|\mathcal{C}^{\Phi}_{0}(n^{\Phi}_{l})|)$ steps.
@@
(Execution time of anti-join nodes is linear in the number of matches for the semi-join’s left dependency.) Let $H$ be a graph, $H_{p}\subseteq H$, $[\rhd]^{\Phi}\in V^{N^{\Phi}}$ a marking-sensitive anti-join node with left dependency $n^{\Phi}_{l}$ and right dependency $n^{\Phi}_{r}$, and $\mathcal{C}^{\Phi}_{0}$ a configuration such that $\mathcal{C}^{\Phi}_{0}([\ltimes]^{\Phi})=\{\}$. Executing $[\ltimes]^{\Phi}$ via $\mathcal{C}^{\Phi}_{1}=execute([\ltimes]^{\Phi},N^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{0})$ then takes $O(|\mathcal{C}^{\Phi}_{0}(n^{\Phi}_{l})|)$ steps.
@@
(RETE net localization via $localize^{\Psi}$ introduces an overhead of at most factor 10 on the number of computed matches.) Let $H$ be an edge-dominated graph, $H_{p}\subseteq H$, $(N,p)$ a RETE net created via the procedure described in [BG23] for the extended graph query $(Q,\psi)$, $\mathcal{C}$ a consistent configuration for $(N,p)$ for host graph $H$, and $\mathcal{C}^{\Phi}$ a consistent configuration for the marking-sensitive RETE net $(N^{\Phi},p^{\Phi})=localize^{\Psi}(Q,\psi)$ for host graph $H$ and relevant subgraph $H_{p}$ corresponding to $(N,p)$. It then holds that $\sum_{n^{\Phi}\in V^{N^{\Phi}}}|\mathcal{C}^{\Phi}(n^{\Phi})|\leq 10\cdot\sum_{n\in V^{N}}\mathcal{C}(n)|$.
@@
In the base case where $\psi=true$, it holds that $localize^{\Psi}(Q,\psi)=localize(N,p)$. From Lemma 26, it then immediately follows that $\sum_{n^{\Phi}\in V^{N^{\Phi}}}|\mathcal{C}^{\Phi}(n^{\Phi})|\leq 10\cdot\sum_{n\in V^{N}}|\mathcal{C}(n)|$.
@@
For a nested condition of the form $\psi=\exists(a:Q\rightarrow Q^{\prime},\psi^{\prime})$, $N^{\Phi}$ consists of the localized RETE net for $Q$, $(N^{\Phi}_{Q},p^{\Phi}_{Q})=localize(Q)$, the localized RETE net $(N^{\Phi}_{(Q^{\prime},\psi^{\prime})},p^{\Phi}_{(Q^{\prime},\psi^{\prime})})=localize^{\Psi}(Q^{\prime},\psi^{\prime})$, a request projection structure $RPS^{\infty}_{l}=RPS^{\infty}(p^{\Phi}_{Q},N^{\Phi}_{(Q^{\prime},\psi^{\prime})})$, and a marking-sensitive semi-join node $p^{\Phi}=[\ltimes]^{\Phi}$.
@@
In the worst case, the extension points of $(N^{\Phi}_{(Q^{\prime},\psi^{\prime})},p^{\Phi}_{(Q^{\prime},\psi^{\prime})})$ already contain all possible single-vertex matches into $H$ even without the additional input via $RPS^{\infty}_{l}$. Therefore, by the induction hypothesis, it must hold that $\sum_{n^{\Phi}\in V^{N^{\Phi}_{(Q^{\prime},\psi^{\prime})}}}|\mathcal{C}^{\Phi}(n^{\Phi})|\leq 10\cdot\sum_{n\in V^{N^{\prime}}}|\mathcal{C}(n)|$.
From Lemma 12 in [BG24b], it follows that $\sum_{n^{\Phi}_{Q}\in V^{RPS^{\infty}_{l}}}|\mathcal{C}^{\Phi}(n^{\Phi}_{Q})|\leq 3\cdot|\mathcal{C}(p_{Q})|$.
@@
Consequently, it must hold that $\sum_{n^{\Phi}\in V^{N^{\Phi}}}|\mathcal{C}^{\Phi}(n^{\Phi})|\leq 10\cdot\sum_{n\in V^{N}}|\mathcal{C}(n)|$.
For a nested condition of the form $\psi=\neg\psi^{\prime}$, $N^{\Phi}$ consists of the localized RETE net for the plain pattern $Q$, $(N^{\Phi}_{Q},p^{\Phi}_{Q})=localize(Q)$, the localized RETE net $(N^{\Phi}_{(Q,\psi^{\prime})},p^{\Phi}_{(Q,\psi^{\prime})})=localize^{\Psi}(Q,\psi^{\prime})$, a request projection structure $RPS^{\infty}_{l}=RPS(p^{\Phi}_{Q},N^{\Phi}_{(Q,\psi^{\prime})})$, and a marking-sensitive anti-join node $p^{\Phi}=[\rhd]^{\Phi}$.
@@
By analogous argumentation as for the existential case, it must then also hold that $\sum_{n^{\Phi}\in V^{N^{\Phi}}}|\mathcal{C}^{\Phi}(n^{\Phi})|\leq 10\cdot\sum_{n\in V^{N}}|\mathcal{C}(n)|$.
For a nested condition of the form $\psi=\psi_{1}\wedge\psi_{2}$, $N^{\Phi}$ consists of the localized RETE net for the plain pattern $Q$, $(N^{\Phi}_{Q},p^{\Phi}_{Q})=localize(Q)$, the localized RETE net $(N^{\Phi}_{(Q,\psi_{1})},p^{\Phi}_{(Q,\psi_{1})})=localize^{\Psi}(Q,\psi_{1})$, a request projection structure $RPS^{\infty}_{1}=RPS^{\infty}(p^{\Phi}_{Q},N^{\Phi}_{(Q,\psi_{1})})$, a marking-sensitive semi-join node $[\ltimes]^{\Phi}_{1}$, the localized RETE net $(N^{\Phi}_{(Q,\psi_{2})},p^{\Phi}_{(Q,\psi_{2})})=localize^{\Psi}(Q,\psi_{2})$, a request projection structure $RPS^{\infty}_{2}=RPS^{\infty}([\ltimes]^{\Phi}_{1},N^{\Phi}_{(Q,\psi_{2})})$, and a marking-sensitive semi-join node $p^{\Phi}=[\ltimes]^{\Phi}_{2}$.
@@
As in the existential case, it holds that $\sum_{n^{\Phi}\in V^{N^{\Phi}_{Q}}}|\mathcal{C}^{\Phi}(n^{\Phi})|\leq 10\sum_{n\in V^{N_{Q}}\setminus\{p_{Q}\}}|\mathcal{C}(n)|+7\cdot|\mathcal{C}(p_{Q})|$. Furthermore, it must hold that $\sum_{n^{\Phi}\in V^{N^{\Phi}_{(Q,\psi_{1})}}}|\mathcal{C}^{\Phi}(n^{\Phi})|\leq 10\cdot\sum_{n\in V^{N_{1}}}|\mathcal{C}(n)|$ and $\sum_{n^{\Phi}\in V^{N^{\Phi}_{(Q,\psi_{1})}}}|\mathcal{C}^{\Phi}(n^{\Phi})|\leq 10\cdot\sum_{n\in V^{N_{2}}}|\mathcal{C}(n)|$ by the induction hypothesis.
From Lemma 12 in [BG24b], it follows that $\sum_{n^{\Phi}_{Q}\in V^{RPS^{\infty}_{1}}}|\mathcal{C}^{\Phi}(n^{\Phi}_{Q})|\leq 3\cdot|\mathcal{C}(p_{Q})|$.
We furthermore know by Theorem 8 that $\mathcal{C}^{\Phi}([\ltimes]^{\Phi}_{1})$ can only contain matches for $Q$ that satisfy $\phi_{1}$ and $\mathcal{C}([\ltimes]_{1})$ contains all such matches. From Lemma 12 in [BG24b], it thus follows that $|\mathcal{C}^{\Phi}([\ltimes]^{\Phi}_{1})|\leq|\mathcal{C}([\ltimes]_{1})|$. Consequently, it must hold that $\sum_{n^{\Phi}_{Q}\in V^{RPS^{\infty}_{2}}}|\mathcal{C}^{\Phi}(n^{\Phi}_{Q})|\leq 3\cdot|\mathcal{C}([\ltimes]_{1})|$. By Theorem 8 and Lemma 12 in [BG24b], it also follows that $|\mathcal{C}^{\Phi}([\ltimes]^{\Phi}_{2})|\leq|\mathcal{C}([\ltimes]_{2})|$.
We thus know that $\sum_{n^{\Phi}\in V^{N^{\Phi}}}|\mathcal{C}^{\Phi}(n^{\Phi})|\leq 10\cdot\sum_{n\in V^{N}}|\mathcal{C}(n)|$.
@@
(Consistent configurations for modular extensions of RETE nets localized via $localize^{sat}$ yield all subgraph satisfaction dependent matches for an extended graph query with marking $\infty$.) Let $H$ be a graph, $H_{p}\subseteq H$, $(Q,\psi)$ an extended graph query, and $\mathcal{C}^{\Phi}$ a consistent configuration for the modular extension $(X^{sat},p^{sat})$ of $(N^{sat},p^{sat})=localize^{sat}(Q,\psi)$. It then holds that $\forall m\in\mathcal{M}^{Q}_{H}:m\text{ is subgraph satisfaction dependent }\Rightarrow(m,\infty)\in\mathcal{C}^{\Phi}(p^{sat})$.
@@
For a nested condition of the form $\psi=\exists(a:Q\rightarrow Q^{\prime},\psi^{\prime})$, $(N^{sat},p^{sat})=localize^{sat}(Q,\psi)$ consists of the localized RETE net for the plain pattern $Q$, $(N^{\Phi}_{Q},p^{\Phi}_{Q})=localize(Q)$, the localized RETE net for the plain pattern $Q^{\prime}$, $(N^{\Phi}_{Q^{\prime}},p^{\Phi}_{Q^{\prime}})=localize(Q^{\prime})$, the RETE net $(N^{sat}_{(Q^{\prime},\psi^{\prime})},p^{sat}_{(Q^{\prime},\psi^{\prime})})=localize^{sat}(Q^{\prime},\psi^{\prime})$, a marking-sensitive union node $[\cup]^{\Phi}$ with dependencies $p^{\Phi}_{Q^{\prime}}$ and $p^{sat}_{(Q^{\prime},\psi^{\prime})}$, a request projection structure $RPS^{\infty}_{r}=RPS^{\infty}([\cup]^{\Phi},N^{\Phi}_{Q})$, and a marking-sensitive semi-join node $p^{sat}=[\ltimes]^{\Phi}$ with left dependency $p^{\Phi}_{Q}$ and right dependency $[\cup]^{\Phi}$.
By Lemma 1 in [BG24b], the induction hypothesis, and the semantics definition of the marking-sensitive union node, it must then hold that $\forall m^{\prime}\in\mathcal{M}^{Q^{\prime}}_{H}:m^{\prime}(Q^{\prime})\cap H_{p}\neq\emptyset\text{ or }m^{\prime}\text{ is subgraph satisfaction dependent }\Rightarrow(m^{\prime},\infty)\in\mathcal{C}^{\Phi}([\cup]^{\Phi})$.
From the construction of $RPS^{\infty}_{r}$ and Lemma 1 in [BG24b], it follows that $\forall m\in\mathcal{M}^{Q}_{H}:(\exists m^{\prime}\in\mathcal{M}^{Q^{\prime}}_{H}:m=m^{\prime}\circ a\wedge(m^{\prime}(Q^{\prime})\cap H_{p}\neq\emptyset\text{ or }m^{\prime}\text{ is subgraph satisfaction dependent}))\Rightarrow(m,\infty)\in\mathcal{C}^{\Phi}(p^{\Phi}_{Q})$.
@@
(Execution of modular extensions of localized RETE nets via $order^{sat}$ yields consistent configurations for the base RETE net.) Let $H$ be a graph, $H_{p}\subseteq H$, $(Q,\psi)$ a graph query, and $(X^{sat},p_{X}^{sat})$ a modular extension of the marking-sensitive RETE net $(N^{sat},p^{sat})=localize^{sat}(Q,\psi)$. Furthermore, let $\mathcal{C}^{\Phi}_{0}$ be an arbitrary configuration for $(X^{sat},p_{X}^{sat})$. Executing $(X^{sat},p_{X}^{sat})$ via $O=order^{sat}(N^{sat})$ then yields a configuration $\mathcal{C}^{\Phi}=execute(O,N^{sat},H,H_{p},\mathcal{C}^{\Phi}_{0})$ that is consistent for all nodes in $N^{sat}$.